Suppose a discrete random variable y has the mean E(Y)=5 and V(Y)=49. What is the value of V(-2Y+3)?
Answer #1

Now , V(-2Y+3)=V(-2Y)+V(3)=(-2)2V(Y)+0=4*49=196

Since , the variance of constant is 0. i.e. V(3)=0 and V(aY)=a2V(Y)